OpenStudy (anonymous):

How does one solve a " literal equation" Ex: Solve S = 4v2 for v

OpenStudy (amistre64):

by using properties of algebra

OpenStudy (amistre64):

the idea is to get the "v" all alone; so start undoing the stuff that is stuck to it

OpenStudy (btaylor):

first, divide both sides by 4. you get:\[\frac{S}{4} = v^2\]Now take the square root of both sides, and you get:\[\pm \frac{\sqrt{S}}{2}=v\]